# Retrieval-Oriented Knowledge for Click-Through Rate Prediction

### v1

@@ 4.3.1. Retrieval Imitation
This knowledge imitation approach serves as a knowledge distillation process, where we extract and inject the retrieval-enhanced knowledge from the pre-trained model into the learnable knowledge base, enabling the implicit sample-level retrieval capability.
@@ 4.3.2. Contrastive Regularization
(4) $\vspace*{-0.5\baselineskip}\mathbf{p}^{x}_{t}$ $\triangleq h\left(\mathbf{z}^{x}_{t}\right)\triangleq h\left(g\left(f\left(x_{t}\right)\right)\right),$
@@
(6) $\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)=\frac{\mathbf{p}^{x}_{t}}{\left\|\mathbf{p}^{x}_{t}\right\|_{2}}\cdot\frac{\mathbf{z}^{s}_{t}}{\left\|\mathbf{z}^{s}_{t}\right\|_{2}},\vspace*{-0.5\baselineskip}$
@@
(7) $\mathcal{L}_{contra}=-\left(\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)+\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{s}_{t},\mathbf{z}^{x}_{t}\right)\right).\vspace*{-0.5\baselineskip}$
@@
(8) $\mathcal{L}=(1-\alpha)\cdot\mathcal{L}_{imit}+\alpha\cdot\mathcal{L}_{contra},\vspace*{-0.5\baselineskip}$
@@ 4.4. Knowledge Utilization
  • Feature-wise retrieval-enhanced method leverages the retrieval-oriented embedding layer $f$ to obtain the retrieval-enhanced feature-wise embeddings for the target sample. Specifically, this method integrates the retrieval-oriented embedding layer $f$ with the original learnable feature embedding of the backbone model to form the final embedding layer. 333 The embedding layer accounts for a large proportion of the total number of parameters in CTR models. For a fair comparison in model parameters, the embedding size of both the retrieval-oriented embedding layer and the original learnable feature embedding of the backbone model is set to $d/2$ compared with other models. .
@@
    $\left\{\begin{array}[]{ll}\mathbf{z}_{t}^{x}&=g(f(x_{t}))\\ \hat{y}_{t}&=\text{backbone}(\phi(x_{t})\oplus\mathbf{z}_{t}^{x})\end{array},\right.$
@@ 5. Experiments
  • RQ1: How does ROK’s performance compare to that of traditional CTR models and retrieval-based methods?
@@
  • RQ4: How does the chosen update strategy and positive sample selecting strategy affect ROK’s performance?
  • RQ5: How do the specified hyperparameter influence ROK’s performance?
@@ 5.1. Experimental Settings
The evaluations were conducted on three widely-recognized public datasets: Tmall, Taobao, and Alipay. Comprehensive statistics for these datasets are presented in Table 1. Specifically, we tallied the counts of users, items, samples, fields, and categories for each dataset. It’s noteworthy to mention that the number of features refers to the count of unique feature values.
@@
Following RIM (Qin et al., 2021), we organized the data such that the oldest instances constituted the search pool, the most recent instances comprised the test set, and those in between were assigned to the training set. The retrieval-based methods (Qin et al., 2021, 2020) retrieve neighboring samples from the search pool. For user behavior modeling methods (Zhou et al., 2018, 2019) involved (e.g., DIEN), the sequential features (e.g., user behavior) are also generated from the search pool.
@@
To measure the performance, we utilize the commonly adopted metrics AUC and log-loss, which reflect pairwise ranking performance and point-wise likelihood, respectively. A significance test contrasting the two top-performing methods for each metric is undertaken, with results of significance denoted by an asterisk (*).
We follow (Zhou et al., 2018) to introduce Rel. Impr. metric to quantify the relative improvement of models, which is defined as below:
@@
For baselines, we compare ROK with a mid-tier performance backbone model against traditional models and retrieval-based models. For traditional CTR models, GBDT (Chen et al., 1996) is a frequently utilized tree-based model, and DeepFM (Guo et al., 2017) is a popular deep-learning model that focuses on feature interactions. HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019) utilize memory network architectures, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are attention-based CTR models designed to pinpoint user interests. Additionally, FATE (Wu et al., 2021) is a model for learning representations of tabular data which facilitates interactions among samples within a minibatch. In item-level retrieval-based models, SIM (Pi et al., 2020) and UBR (Qin et al., 2020) extract pertinent user behaviors from a comprehensive set of user-generated data. As for the sample-level retrieval-based model, RIM (Qin et al., 2021) retrieves relevant data instances based on raw input features, and DERT (Zheng et al., 2023) retrieves dense representations. To ensure a fair comparison, we adopted the same hyper-parameter settings for the baselines and the ROK backbone during the knowledge utilization phase as those used in prior work (Qin et al., 2021).
Regarding backbone models, we intentionally select average-performing common models like DIN and DIEN rather than top-tier ones like UBR, which will further improve the performance of ROK. By default, we choose DIEN (Zhou et al., 2019) for Tmall and Taobao datasets and DIN (Zhou et al., 2018) for the Alipay dataset. The substantial improvement observed in 5.2 when these backbone models are combined with ROK underscores the efficacy of ROK.
@@ 5.2. Overall Performance Comparison: RQ1
In this section, we compare the overall experimental results in Table 2, from which we have the following observations. First, ROK remarkably boosts the backbone model with AUC improved by 10.08%, 34.96%, and 16.85% on the three datasets respectively, demonstrating the superior performance. Second, retrieval-based methods (Qin et al., 2020, 2021) significantly outperform user behavior modeling methods (Zhou et al., 2018, 2019), benefiting from their superior knowledge retrieval ability. Moreover, sample-level retrieval-based methods (Qin et al., 2021) perform better than item-level retrieval-based methods (Pi et al., 2020; Qin et al., 2020) due to more comprehensive retrieval. Finally, ROK markedly outperforms the item-level retrieval-based methods such as SIM (Pi et al., 2020) and UBR (Qin et al., 2020), and achieves performance surpassing that of the sample-level retrieval-based methods like RIM (Qin et al., 2021). Although ROK is only based on knowledge distillation and contrastive learning, it not only surpasses the teacher model but also comes very close to recent results in terms of performance. Notably, in our methodology, we deliberately opted for models of average performance, such as DIN and DIEN, over elite models like UBR (Qin et al., 2020), with the anticipation that the latter would enhance the efficacy of ROK. Concurrently, it is pertinent to note that the teacher model employed within ROK is RIM, which DERT outperforms. Nonetheless, the adoption of a superior teacher model is poised to markedly augment the performance of ROK. Despite this, ROK has successfully enabled user behavior modeling methods to outperform both item-level retrieval-based methods and sample-level retrieval-based methods. This indicates that the neural knowledge model successfully serves as a compact surrogate for the retrieval pool. More importantly, the sample-level retrieval-based methods, due to their prolonged inference time, were once deemed entirely infeasible for industrial application, which will further be elaborated in Section 6.2. However, with the introduction of ROK, this barrier has been overcome. ROK not only mitigates the challenges but revolutionizes the deployment prospects, rendering sample-level retrieval-based methods feasible industry applications.
Table 2. Performance comparison of CTR prediction task with various baselines. For each dataset, RIM is selected as the teacher model, and an average-performing backbone model is selected for ROK: DIEN for both the Tmall and Taobao datasets and DIN for the Alipay dataset. indicates recent results. Evaluation metrics include AUC and log-loss (LL). Among the models excluding recent work, the best results are highlighted in bold, while the runner-ups are underlined. The gray-colored bold text indicates recent results outperform all other models. “Rel. Impr.” signifies the model’s relative AUC improvement over the chosen backbone model, with a statistical significance level of $p<0.01$.
@@ 5.3. Compatibility Analysis: RQ2
In this section, we commence by contrasting the compatibility of our novel approach, ROK, with conventional CTR models, with the findings delineated in Table 3. Upon examining the data, two observations arise: (i) The application of ROK markedly amplifies the performance of the backbone models. To illustrate, by integrating ROK with DeepFM, DIN, and DIEN, there’s a notable enhancement in performance metrics, with an average AUC increase of 4.30%, 8.01%, and 5.59%, coupled with an average decrease in log-loss by 11.80%, 8.51%, and 8.22% across three datasets, respectively. This highlights the significant benefits of incorporating the retrieval-oriented knowledge from ROK into the backbone models. (ii) Exhibiting model-agnosticism, ROK seamlessly integrates with a diverse array of backbone CTR models, irrespective of whether they focus on feature interaction or behavior modeling. This underscores that the knowledge garnered by ROK is beneficial across diverse models, highlighting the compatibility of ROK. The feature of model-agnosticism is crucial since, regardless of the optimal model for a given scenario, ROK consistently boosts its capabilities. Thus, instead of being bound to one model, we can choose the ideal one for each situation and leverage ROK to optimize its performance.
@@ 5.4. Comparison of the learned knowledge of ROK and RIM: RQ3
In this section, we delve into the comparative analysis of the knowledge quality acquired by ROK and RIM. For a detailed assessment, we formulated two simplest variants (LR and MLP) for each model to fully demonstrate the effect of knowledge:
@@
  • For ROK, we designed the ROK(LR) and ROK(MLP) variants. Both of these take as input the concatenation of the original feature embeddings $\mathbf{x}_{t}$, with ROK’s knowledge component $z_{t}^{x}$.
@@
  • Correspondingly, for RIM, we introduced the RIM(LR) and RIM(MLP) variants. Their input comprises the concatenation of the original feature embeddings, $\mathbf{x}_{t}$, with RIM’s aggregated features and labels, $\mathbf{r}_{t}$.
@@
The detailed performance comparison of these variants is tabulated in Table 4 and we have two observations. First, significantly, the results indicate that the ROK variants surpass the RIM variants in performance across both datasets. This distinction in outcomes emphasizes the superior quality of knowledge learned by ROK, especially as these models apply this intrinsic knowledge directly for predictions. It’s compelling to note that while ROK leverages the retrieval imitation strategy (detailed in Section 4.3.1) to imitate the aggregated features and labels of RIM, it extracts richer and more profound knowledge from the data. We are inclined to believe that the contrastive regularization methodology is pivotal in driving this enhanced performance for ROK. Second, ROK(MLP) in Tmall even outperforms DeepFM and some user behavior modeling methods, which shows another way to improve the pre-ranking model effect.
@@
To delve deeper into the inherent knowledge of ROK and explain the superiority of ROK over RIM, we employed t-SNE visualization (Van der Maaten and Hinton, 2008). Figure 4 and Figure 7 depict the knowledge distribution patterns of both ROK and RIM across the Alipay and Tmall datasets. For this visualization, we randomly selected 10,000 samples from both the training and testing sets for each model. We have two observations. First, a discerning observation from Figure 4 reveals that ROK’s knowledge distribution exhibits a more pronounced clustering effect across all datasets. This enhanced clustering can be attributed to contrastive regularization. For comparison, RIM’s retrieved neighboring samples may include some incidental noise that correlates with the quantity of these samples. Besides, the aggregation mechanism of RIM only captures the global features of the neighboring samples. However, the contrastive regularization in ROK emphasizes the local features of neighboring samples and mitigates the noise of neighboring samples by selectively considering the most relevant sample, as shown in section 4.3.2. Specifically, representations of analogous samples are pulled closer, while those of dissimilar samples are pushed apart, leading to a more distinct and meaningful clustering. Secondly, a significant concern when using neural networks as a knowledge base in our case relates to their generalization abilities. This is because the training process is guided by the retrieved examples of sample-level retrieval-based methods, while the testing phase requires the model to apply its knowledge independently. Nonetheless, the visualization of ROK’s representations reveals a high degree of similarity between the training and test datasets, showing the good generalization of the decomposition-reconstruction paradigm.
@@ 5.5. Ablation & Hyper-parameter Study
In this section, we undertake comprehensive ablation experiments for ROK. We start by examining the parameter update strategy employed for the knowledge base during the knowledge utilization phase. The strategies under consideration include: fixing the knowledge base (Fix), updating only the upper knowledge encoder $g$ (Upd $g$), and updating both the knowledge encoder and the retrieval-oriented embedding layer $f$ (Upd $f+g$). Furthermore, we consider the positive sample selection strategy mentioned in Section 4.3.2. Specifically, we select the most related neighboring sample by default. Now, we compare the performance of randomly selecting a sample from the retrieved samples (random). Results obtained on the Tmall and Alipay datasets are depicted in Figure 5. We have two observations: First, we observe that randomly selecting a sample as the positive sample diminishes performance. This suggests that even within the narrowly retrieved set of samples, noise is still present. Second, a key observation from these results is that the strategy of fixing the knowledge base (Fix) yields the best performance. This can be attributed to the primary goal of ROK: to derive superior representations. However, updating with the backbone models may adversely impact the acquired knowledge. This characteristic is highly desirable and aligns well with our expectations, further enhancing the practicality of ROK. This indicates that once ROK has been fully trained, it can function as an independent module. This eliminates the need for subsequent modifications, reinforcing its role as an authentic knowledge base.
@@
In this section, we delve into the study of the hyperparameter $\alpha$ pertaining to the balance between the contrastive regularization loss and retrieval imitation loss, represented as $\alpha$. As illustrated in Figure 6, for both datasets, when $\alpha$ assumes values of either 00 or $1$—representing the extreme scenarios—the performance of ROK deteriorates. This observation underscores the significance of both retrieval imitation loss and contrastive regularization loss to the model’s efficacy.
((a)) Tmall

### v2

@@ 4.3.1. Retrieval Imitation
This knowledge imitation approach acts as a knowledge distillation process, where we distill knowledge from a non-parametric retrieval process using a parametric approach. By extracting and injecting the retrieval-oriented knowledge from the pre-trained model into the learnable knowledge base, we enable implicit sample-level retrieval capability.
@@ 4.3.2. Contrastive Regularization
(4) $\mathbf{p}^{x}_{t}$ $\triangleq h\left(\mathbf{z}^{x}_{t}\right)\triangleq h\left(g\left(f\left(x_{t}\right)\right)\right),$
@@
(6) $\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)=\frac{\mathbf{p}^{x}_{t}}{\left\|\mathbf{p}^{x}_{t}\right\|_{2}}\cdot\frac{\mathbf{z}^{s}_{t}}{\left\|\mathbf{z}^{s}_{t}\right\|_{2}},$
@@
(7) $\mathcal{L}_{contra}=-\left(\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)+\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{s}_{t},\mathbf{z}^{x}_{t}\right)\right).$
@@
(8) $\mathcal{L}=(1-\alpha)\cdot\mathcal{L}_{imit}+\alpha\cdot\mathcal{L}_{contra},$
@@ 4.4. Knowledge Utilization
  • Feature-wise retrieval-enhanced method leverages the retrieval-oriented embedding layer $f$ to obtain the retrieval-enhanced feature-wise embeddings for the target sample. Specifically, this method integrates the retrieval-oriented embedding layer $f$ with the original learnable feature embedding of the backbone model to form the final embedding layer444 The embedding layer accounts for a large proportion of the total number of parameters in CTR models. For a fair comparison in model parameters, the embedding size of both the retrieval-oriented embedding layer and the original learnable feature embedding of the backbone model is set to $d/2$ compared with other models. .
@@
    (11) $\left\{\begin{array}[]{ll}\mathbf{z}_{t}^{x}&=g(f(x_{t}))\\ \hat{y}_{t}&=\text{backbone}(\phi(x_{t})\oplus\mathbf{z}_{t}^{x})\end{array},\right.$
@@ 5. Experiments
  • RQ1: How does ROK’s performance compare to that of traditional CTR models and the teacher model?
@@
  • RQ4: How do the chosen update strategy and positive sample selecting strategy affect ROK’s performance?
  • RQ5: How does the specified hyperparameter influence ROK’s performance?
@@ 5.1. Experimental Settings
The evaluations are conducted on three widely-recognized public datasets: Tmall, Taobao, and Alipay. Comprehensive statistics for these datasets are presented in Table 1. To be specific, we tally the counts of users, items, samples, fields, and categories for each dataset. It’s worth noting that the number of features refers to the count of unique feature values.
@@
Following RIM (Qin et al., 2021), we organize the data such that the oldest instances constitute the search pool, the most recent instances comprised the test set, and the instances in between are assigned to the training set. The retrieval-based methods (Qin et al., 2021, 2020) retrieve neighboring samples from the search pool. For user behavior modeling methods (Zhou et al., 2018, 2019) (e.g., DIEN), the sequential features (e.g., user behavior) are also generated from the search pool.
@@
To measure performance, we utilize the commonly adopted metrics AUC and log-loss, which reflect pairwise ranking performance and point-wise likelihood, respectively. A significance test contrasting the two top-performing methods for each metric is undertaken, with results of significance denoted by an asterisk (*).
We follow previous work (Zhou et al., 2018) to introduce Rel. Impr. metric to quantify the relative improvement of models, which is defined as below:
@@
For baselines, we compare ROK with a mid-tier performance backbone model against traditional models and retrieval-based models. For traditional CTR models, GBDT (Chen et al., 1996) is a frequently used tree-based model, and DeepFM (Guo et al., 2017) is a popular deep-learning model that focuses on feature interactions. HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019) utilize memory network architectures, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are attention-based CTR models designed for pinpointing user interests. Additionally, FATE (Wu et al., 2021) is a model for learning representations of tabular data which facilitates interactions among samples within a minibatch. In item-level retrieval-based models, SIM (Pi et al., 2020) and UBR (Qin et al., 2020) extract pertinent user behaviors from a comprehensive set of user-generated data. Regarding the sample-level retrieval-based model, RIM (Qin et al., 2021) retrieves relevant data instances based on raw input features. We do not include DERT (Zheng et al., 2023) in our comparison because it is a special case of RIM that introduces a new dense retrieval setting. Moreover, our primary goal is to develop a method that eliminates the need for retrieval during inference, thereby improving efficiency and enabling real-world deployment. Therefore, we focus more on comparing the performance of ROK with the teacher model i.e., RIM (Qin et al., 2021) to determine whether ROK would cause a significant performance degradation and whether it could enhance the backbone model.
@@ 5.2. Overall Performance Comparison: RQ1
In this section, we compare the overall experimental results presented in Table 2 and make the following observations. First, ROK significantly enhances the performance of the backbone model, with AUC improvements of 10.08%, 34.96%, and 16.85% on the three datasets, respectively, demonstrating its superior performance. Second, retrieval-based methods (Qin et al., 2020, 2021) considerably outperform user behavior modeling methods (Zhou et al., 2018, 2019), benefiting from their advanced knowledge retrieval capabilities. Furthermore, sample-level retrieval-based methods (Qin et al., 2021) exhibit better performance than item-level retrieval-based methods (Pi et al., 2020; Qin et al., 2020) due to more comprehensive retrieval strategies. Finally, ROK notably surpasses the performance of item-level retrieval-based methods methods, such as SIM (Pi et al., 2020) and UBR (Qin et al., 2020), and even outperforms our teacher model, RIM (Qin et al., 2021), which belongs to the sample-level retrieval-based methods. Our primary goal was to develop a method that eliminates the need for retrieval during inference, thereby improving efficiency and enabling real-world deployment. Initially, we expected that our approach would merely approximate the performance of the teacher model while preserving extraordinary inference efficiency, given that it is based on knowledge distillation. However, through the introduction of contrastive regularization, our method not only matches but exceeds the performance of the teacher model. This demonstrates the strong potential of our approach for real-world applications, as it can effectively learn and utilize the knowledge in the search pool without the need to perform real retrieval during inference. Our results highlight the effectiveness of knowledge distillation combined with contrastive regularization in capturing and leveraging the information from the search pool, enabling efficient and high-performing models suitable for deployment in practical scenarios.
Table 2. Performance comparison of CTR prediction task with various baselines. For each dataset, RIM is selected as the teacher model, and an average-performing backbone model is selected for ROK: DIEN for both the Tmall and Taobao datasets and DIN for the Alipay dataset. indicates recent results. Evaluation metrics include AUC and log-loss (LL). The best results are highlighted in bold, while the runner-ups are underlined. “Rel. Impr.” signifies the model’s relative AUC improvement over the chosen backbone model, with a statistical significance level of $p<0.01$.
@@ 5.3. Compatibility Analysis: RQ2
In this section, we begin by contrasting the compatibility of our novel approach, ROK, with conventional CTR models, as delineated in Table 3. Upon examining the data, two observations arise: (i) The application of ROK markedly amplifies the performance of the backbone models. To illustrate, by integrating ROK with DeepFM (Guo et al., 2017), DIN (Zhou et al., 2018), and DIEN (Zhou et al., 2019), there is a notable enhancement in performance metrics, with an average AUC increase of 4.30%, 8.01%, and 5.59%, coupled with an average decrease in log-loss by 11.80%, 8.51%, and 8.22% across three datasets, respectively. This highlights the significant benefits of incorporating the retrieval-oriented knowledge from ROK into the backbone models and indicates that the neural knowledge model successfully serves as a compact surrogate for the search pool. (ii) ROK, exhibiting model-agnosticism, seamlessly integrates with a diverse array of backbone CTR models, irrespective of whether they focus on feature interaction or behavior modeling. This underscores that the knowledge garnered by ROK is beneficial across diverse models, highlighting the compatibility of ROK. The feature of model-agnosticism is crucial since, regardless of the optimal model for a given scenario, ROK consistently boosts its capabilities. Thus, instead of being bound to one model, we can choose the ideal one for each situation and leverage ROK to optimize its performance.
@@ 5.4. Comparison of the learned knowledge of ROK and RIM: RQ3
In this section, we delve into the comparative analysis of the knowledge quality acquired by ROK and RIM. For a detailed assessment, we formulate two simple variants (LR and MLP) for each model to fully demonstrate the effect of knowledge:
@@
  • For ROK, we design the ROK(LR) and ROK(MLP) variants. Both take as input the concatenation of the original feature embeddings $\mathbf{x}_{t}$, with ROK’s knowledge component $\mathbf{z}^{x}_{t}$.
@@
  • Similarly, for RIM (Qin et al., 2021), we introduced the RIM(LR) and RIM(MLP) variants. Their input comprises the concatenation of the original feature embeddings, $\mathbf{x}_{t}$, with RIM’s aggregated features and labels, $\mathbf{r}_{t}$.
@@
The detailed performance comparison of these variants is tabulated in Table 4 and we make two observations. First, significantly, the results indicate that the ROK variants surpass the RIM variants in performance across both datasets. This distinction in outcomes emphasizes the superior quality of knowledge learned by ROK, especially as these models apply this intrinsic knowledge directly for predictions. It is compelling to note that while ROK leverages the retrieval imitation strategy (detailed in Section 4.3.1) to imitate the aggregated features and labels of RIM, it extracts richer and more profound knowledge from the data. We believe that the contrastive regularization methodology is pivotal in driving this enhanced performance for ROK. Second, ROK(MLP) in Tmall even outperforms DeepFM (Guo et al., 2017) and some user behavior modeling methods, which shows another way to improve the pre-ranking model effect.
@@
To delve deeper into the inherent knowledge of ROK and explain its superiority over RIM, we employed t-SNE visualization (Van der Maaten and Hinton, 2008). Figure 4 illustrates the knowledge distribution patterns of both ROK and RIM in the Alipay dataset. For this visualization, we randomly selected $10,000$ samples from both the training and testing sets for each model. We have two observations. First, ROK’s knowledge distribution exhibits a more pronounced clustering effect compared to RIM. This enhanced clustering can be attributed to the contrastive regularization employed by ROK. The contrastive regularization emphasizes the local features of neighboring samples and mitigates the noise by selectively considering the most relevant sample, as shown in section 4.3.2. Specifically, representations of analogous samples are pulled closer, while those of dissimilar samples are pushed apart, leading to a more distinct and meaningful clustering. In contrast, RIM’s retrieved neighboring samples may include some incidental noise that correlates with the quantity of these samples, and its aggregation mechanism only captures the global features of the neighboring samples. Second, the visualization of ROK’s representations reveals a high degree of similarity between the training and test datasets, demonstrating the good generalization ability of the decomposition-reconstruction paradigm. This is significant because a major concern when using neural networks as a knowledge base is their generalization ability, as the training process is guided by the retrieved examples of sample-level retrieval-based methods, while the testing phase requires the model to apply its knowledge independently.
@@ 5.5. Ablation & Hyper-parameter Study
In this section, we undertake comprehensive ablation experiments for ROK. We begin by examining the parameter update strategy employed for the knowledge base during the knowledge utilization phase. The strategies under consideration include: fixing the knowledge base (Fix), updating only the upper knowledge encoder $g$ (Upd $g$), and updating both the knowledge encoder and the retrieval-oriented embedding layer $f$ (Upd $f+g$). Furthermore, we consider the positive sample selection strategy described in Section 4.3.2. Specifically, we select the most related neighboring sample by default and compare its performance to randomly selecting a sample from the retrieved samples (random). Figure 5 depicts the results obtained on the Tmall, Taobao and Alipay datasets. We have two observations: First, randomly selecting a sample as the positive sample diminishes performance, suggesting that even within the small retrieved set of samples, noise is still present. Second, a key observation is that the strategy of fixing the knowledge base (Fix) yields the best performance. This can be attributed to the primary goal of ROK: to derive superior representations. However, updating with the backbone models may adversely impact the acquired knowledge. This characteristic is highly desirable and aligns well with our expectations, further enhancing the practicality of ROK. It indicates that once ROK has been fully trained, it can function as an independent module, eliminating the need for subsequent modifications and reinforcing its role as an authentic knowledge base.
@@
In this section, we delve into the study of hyperparameter $\alpha$ pertaining to the balance between the contrastive regularization loss and retrieval imitation loss, represented as $\alpha$. As illustrated in Figure 6, for both datasets, when $\alpha$ equals either 00 or $1$—representing the extreme scenarios—the performance of ROK deteriorates. This observation underscores the significance of both retrieval imitation loss and contrastive regularization loss to the model’s efficacy.
((a)) Taobao